# Jing Liu EnQi Lian

### v1

@@
Abstract Sound Fusion with Unconditioned Inversion Model
Jing Liu EnQi Lian
@@ Abstract
Inversion models, sound fusion, diffusion models, music generation
@@ 1 Introduction
Sound fusion seeks to combine an original sound with a reference sound, creating a new auditory output whose acoustic characteristics transcend simple additive mixing of the input components. In this study, sounds are classified into two primary categories: concrete sounds and abstract sounds. Concrete sounds are those that listeners can clearly associate with specific objects, scenes, events, or musical patterns that are intuitively recognizable due to their semantic content. Examples such as running water, crying children, arriving trains, metallic percussion, and burning firewood can be readily identified as concrete. Conversely, abstract sounds are those that do not reveal identifiable real-world sound events or, in musical contexts—particularly in contemporary, electronic, and experimental music clips—do not intuitively disclose recognizable patterns. In essence, the presence of describable, semantically meaningful patterns determines whether a sound is classified as concrete or abstract.
@@
This research specifically focuses on the fusion of abstract sounds. While concrete sounds offer limited potential for meaningful fusion due to their inherent semantic specificity, fused abstract sounds demonstrate musical potential that serve as material for composition.
While we use ”sound” to distinguish our focus from music, this research technically operates on digital audio as the carrier medium.
The main contributions of this work are as follows:
@@
  • We introduce a sound fusion approach that synthesizes auditory features from an original sound and a reference sound.
  • We propose both Stochastic Differential Equation (SDE) and Ordinary Differential Equation (ODE) inversion models based on DPMSolver++ samplers that enable this sound fusion capability.
@@ 2 Related Works
Audio editing has evolved from earlier approaches that primarily focused on applications such as voice conversion, singing voice conversion, and timbre transfer, with research concentrated in speech and music domains. The emergence of Text-To-Audio (TTA) systems in recent years has
$\boldsymbol{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{\boldsymbol{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon^{(t)}_{\theta}}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\epsilon^{(t)}_{\theta}+\sigma_{t}\boldsymbol{\epsilon}_{t}$ (1)
broadened the scope of real-sample editing and enhanced precision through prompt manipulation.
These TTA systems typically consist of integrated pipelines comprising pre-trained models, offering diverse entry points for addressing editing challenges. Models such as CLAP (Wu* et al., 2023) and T5 (Raffel et al., 2023), trained to predict samples from text prompts, form the foundation of TTA systems and enable prompt-based editing capabilities.
@@
Recent research in real audio editing has drawn inspiration from image editing approaches (Mokady et al., 2022; Huberman-Spiegelglas et al., 2024; Roich et al., 2022), introducing new editing tasks such as substituting sound events within audio, as well as adding, removing, replacing, and refining concrete audio content (Wang et al., 2023; Xu et al., 2024). Furthermore, the research scope has expanded from music and speech to encompass general audio.
@@
Most pre-existing inversion studies have aimed to reverse samples back to their initial latent codes $x_{T}$, from which sampling begins. This technique originated during the GAN era (Zhu et al., 2018; Xia et al., 2022; Meng et al., 2022) and facilitates various applications, including in-painting and super-resolution. From an editing perspective, inversion addresses the challenges of unexpected changes and unwanted feature preservation that occur when employing sample-to-sample approaches. Research in this area has predominantly focused on computer vision applications, with the audio field subsequently adapting these techniques, though for both domains, inversion research on diffusion models remains relatively limited.
@@
Prior inversion approaches have worked with DDPM (Ho et al., 2020) and DDIM (Song et al., 2022), either approximating inversion through diffusion itself or by reversing the sampling process. (Huberman-Spiegelglas et al., 2024) computed latent codes at subsequent timesteps directly from $x_{0}$, and recent work (Manor & Michaeli, 2024) extended this approach to the audio domain. (Mokady et al., 2022) worked with the DDIM sampling ODE and swapped timesteps for inversion approximation. Within the context of Classifier-Free Guidance (CFG) (Ho & Salimans, 2022), the authors optimized the negative prompt during inversion while preserving the flexibility of the positive prompt. This approach fine-tuned interpolations against diffusion latents as pivots (Roich et al., 2022). However, fine-tuning embeddings for each edit incurs relatively high computational costs. PPAE (Xu et al., 2024) subsequently utilized this approach for audio editing.
We choose the open-source TTA system Stable Audio Open 1.0 (Evans et al., 2024b) to achieve the best possible perceptual quality. Since this system uses DPMSolver++ (Lu et al., 2023) samplers for which no inversion methods have been proposed, we introduce DPMSolver++ ODE and SDE inversion models to fill this gap. Our approach preserves prompt flexibility, including negative prompts during sampling. Negative prompts, previously termed unconditioned prompts, were originally used as internal mechanisms to introduce variety, but now provide opposing guidance directions in many modern generation models (Rombach et al., 2021; Evans et al., 2024a). We evaluate our inversion latents regarding identity preservation and prompt responsiveness, and empirical evidence demonstrates that latents computed by our models achieve a balance between both objectives.
We first review the sampling process of a diffusion model in 1. It yields a sample at the previous timestep $x_{t-1}$ from $x_{t}$, where $\sigma_{t}=\eta\sqrt{\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}}\sqrt{1-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}$. Setting $\eta=0$ results in a deterministic DDIM scheme. Setting $\eta\in(0,1]$ results in a stochastic DDIM scheme that interpolates between deterministic sampling and DDPM. When $\eta=1$, the stochastic DDIM is equivalent to DDPM.
2.1 DDIM Inversion
(Mokady et al., 2022) proposed an inversion model based on deterministic DDIM. The authors omitted the stochastic term from Eq. 1 and reversed the timestep direction for inversion. The latent code of the next timestep is then computed as:
$\boldsymbol{x}_{t}:=\sqrt{\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}\boldsymbol{x}_{t-1}+\left(\sqrt{\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t}}}-\sqrt{\frac{1-\bar{\alpha}_{t-1}}{\bar{\alpha}_{t-1}}}\right)\epsilon^{(t-1)}_{\theta}$
They further fine-tuned the inverted latents by computing the loss against latents produced from the diffusion process, optimizing negative prompt embeddings initialized to null, $\psi(^{\prime\prime})$.
2.2 DDPM Inversion
Huberman-Spiegelglas et al. (2024) claimed that continuous noise vectors $\boldsymbol{\epsilon}_{t}$ were tightly correlated, which led to unpredictable manipulation for image editing. They found that computing directly from $\boldsymbol{x}_{0}$ during the forward process imprinted the original signal better on latents. They first computed latents at each timestep with a diffusion process, and altered the noise accumulated from timestep $1$ to $t$ to independent Gaussian noise:
$\boldsymbol{x}_{t}=\frac{\sigma_{t}}{\sigma_{s}}e^{-(\lambda_{t}-\lambda_{s})}\boldsymbol{x}_{s}+2\alpha_{t}\int_{\lambda_{s}}^{\lambda_{t}}e^{-2(\lambda_{t}-\lambda)}\boldsymbol{x}_{\theta}^{(\lambda)}\text{d}\lambda+\sqrt{2}\sigma_{t}\int_{\lambda_{s}}^{\lambda_{t}}e^{-(\lambda_{t}-\lambda)}\text{d}\omega_{\lambda}$ (2)
$\boldsymbol{x}_{r}=\frac{2r_{1}(\sigma_{r}^{2}+1)}{\alpha_{t}(1-e^{-2h})}\left(\boldsymbol{x}_{t}-\frac{\sigma_{t}}{\sigma_{s}}e^{-h}\boldsymbol{x}_{s}+\frac{\alpha_{t}(1-2r_{1})(1-e^{-2h})}{2r_{1}}\boldsymbol{x}_{\theta}^{(s)}+\frac{\sigma_{r}\alpha_{t}(1-e^{-2h})}{2r_{1}\sqrt{\sigma_{r}^{2}+1}}\epsilon^{(r)}_{\theta}\right)$ (3)
$\boldsymbol{x}_{t}=\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t},\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(0,I)$ (4)
Next, they rearranged Eq. 4 to solve for the noise map $\boldsymbol{\epsilon}_{t}\sim(\boldsymbol{x}_{t},\epsilon^{(t)}_{\theta}(\boldsymbol{x}_{t},\psi))$ at each timestep during sampling, given the diffused latents. To reduce accumulated error, the authors then updated $\boldsymbol{x}_{t-1}$ through the DDPM sampling equation.
@@ 3 Methods
In this section, we first introduce the notation employed throughout this paper. We then demonstrate that the noise prediction term introduces circular dependency, leading to numerically unstable fine-tuning. Subsequently, we propose inversion models that reconstruct the original audio and fine-tune Stable Audio Open to reconstruct reference audio. Finally, we present the audio fusion algorithm.
@@ 3.1 Notation and Definitions
In DDIM and DDPM, $\alpha$ represents the portion of the signal, and $\bar{\alpha}$ is the cumulative product of $\alpha$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$. The notation for $\alpha$ and $\bar{\alpha}$ varies in the original papers; however, we use them consistently throughout this paper.
In DPMSolver++, regarding the notation of timesteps $t$ and $s$, $t$ falls within the range $[0,s]$ for $s>0$. Specifically, $t$ precedes $s$ temporally. Lu et al. (2022) transformed the diffusion ODE from the time domain to the half log-SNR domain, which is represented by $\lambda$ and defined as $\lambda:=\log(\frac{\alpha_{t}}{\sigma_{t}})$.
@@
3.2 Predicted Noise Introduces Circular Dependency
Lu et al. (2023) proved that many high-order samplers are unstable at large guided scales, which amplified the derivatives of the noise prediction model and affect the convergence range of the ODE solver. The authors proposed replacing the noise prediction term with the model output $x_{\theta}$ to achieve more stable convergence.
Taking a general SDE sampler as an example, given $s>0$, the latent at the previous timestep $t\in[0,s]$ is denoted as shown in Eq. 2.
Previous research performed inversion on predicted noise. However, expanding the model output with the predicted noise expression results in circular dependency. To address this circular dependency issue intuitively, we substitute the general formula in Eq. 2 into the second-order SDE case and target the latent at the next timestep $r$, as shown in Eq. 3.
From this example, $\epsilon_{\theta}$ takes $\boldsymbol{x}_{r}$ as its input, which is the variable we are solving for on the left side.
@@ 3.3 Model Output as Constant
DPMSolver++ introduced model output to replace the predicted noise term, which provides us with a sampling perspective: for every step to the previous timestep, we desire $\epsilon_{\theta}$ to predict a noise map that perfectly matches the latent such that, given both conditions, the model output is exactly the same as our target sample for inversion. Hence, by assigning the encoded $\boldsymbol{x}_{0}$ of the real sample to $x_{\theta}$ in Eq. 2 and removing the stochastic term, the latent at the next timestep $\boldsymbol{x}_{t}$ is given by:
@@
Here, we have inversion models for both SDE and ODE variants of DPMSolver++. We can further derive the noise map at $t$ from Eq. 5:
$\boldsymbol{\epsilon}^{(t)}=\frac{\boldsymbol{x}_{t}-\alpha_{t}\boldsymbol{x}_{0}}{\sigma_{t}}$ (9)
@@ 3.4 Audio Fusion Algorithm
The fusion process is a modified sampling process. The algorithm operates in two phases: in the first half of fusion, the reference audio prompt dominates to produce interpolations with implicit features of the reference audio. In the second half, we inject the inverted latent into the sampler.
Each term holds explicit significance at the moment of injection. As shown in Eq. 2, the first term is where the inverted interpolation is injected, while the second term holds the interpolations guided by the reference prompt. Collectively, these two terms represent an addition of two audio sources, but the third term provides flexibility for fusion.
The algorithm uses low-order sampling at the beginning of the sampling process to warm up for higher-order sampling, and employs first-order sampling near the end of the sampling process to reduce the impact of numerical instability. Algorithm 1 demonstrate second-order sampling as an example. We ensure that the first- and second-order inversion and sampling remain consistent across timesteps. Additionally, we incorporate the CFG mechanism into the editing process. Let $\boldsymbol{x}$ denote the latent at the current timestep and $\boldsymbol{x}_{i-1}$ denote the latent at the previous timestep.
Algorithm 1 Audio Fusion
  Input: prompt of the reference audio ${}^{\prime}C^{\prime}$, negative prompt ${}^{\prime}U^{\prime}$, inverted latent space $\{\hat{x}\}_{\text{org}}^{i=1,...T}$, injection timestep $t_{j}$, CFG guidance scale $\omega$, warm-up timestep $n_{s}$, cool-down timestep $n_{f}$;
  Output: fused audio $\mathcal{B}$.
  Known: noise prediction network $\epsilon_{\theta}$, noise $z$.
  Project ${}^{\prime}C^{\prime}$ to the embedding $\psi(^{\prime}C^{\prime})$;
  Project ${}^{\prime}U^{\prime}$ to the embedding $\psi(^{\prime}U^{\prime})$;
  Initialize latent $x$.
  for $i=T,...,1$ do
     if $i==t_{j}$ then
        Assign $x\leftarrow\hat{x}_{i}$
     else
        Assign $x\leftarrow x_{i}$
     end if
     Compute the noise residue $\epsilon_{\theta}(x,\psi(^{\prime}C^{\prime}))$;
     Compute the noise residue $\epsilon_{\theta}(x,\psi(^{\prime}U^{\prime}))$;
     Compute the CFG noise residue $\epsilon_{\theta}(x)=\omega\epsilon_{\theta}(x,\psi(^{\prime}C^{\prime}))+(1-\omega)\epsilon_{\theta}(x,\psi(^{\prime}U^{\prime}))$;
     Compute $x_{\theta}(x,\epsilon_{\theta}(x))$ with Eq. 5.
     if $i\geq n_{s}$ or $i<n_{f}$ then
        Compute the first order $x_{i-1}$ using Eq. 2;
     else
        Compute the second order $x_{i-1}$ using Eq. 2.
     end if
     Assign $x_{i}\leftarrow x$.
  end for
  Decode $x_{i-1}$ with the VAE decoder to obtain the fused audio $\mathcal{B}$.
  return $\mathcal{B}$
@@ 4 Experiments and Results
The potential for an inverted latent to reconstruct the original sample must be measured perceptually. Metrics evaluating distance or similarity between vectors, including MSE and cosine similarity, do not necessarily correlate with auditory perceptual similarity, which is our primary concern. While perceptual measurements exist for speech, these metrics are tailored to speech-specific characteristics and do not translate well to general audio. Significant differences exist between speech and general audio regarding frequency bands, harmonic content, and spectral characteristics. Most notably, speech typically occupies a narrower frequency range, whereas general audio spans the full audible spectrum and includes diverse timbral qualities, complex textures, and non-harmonic elements that speech-based perceptual metrics cannot adequately capture. To the best of our knowledge, established metrics for measuring perceptual similarity in general audio are lacking. Therefore, we employ spectrograms to provide intuitive visual analysis in lieu of numerical assessment. [^1] All spectrograms presented in this paper are generated using Sonic Visualiser software. The corresponding audio is available online. [^2] https://github.com/Jng-L/SoundFusion
(a) Employ inverted latents from our method.
(b) Employ latents with noise added from forward diffusion.
(c) Employ noisy latents with stochastic terms preserved from our method.
Figure 2: Second-order SDE reconstruction with null prompt guidance. Columns 1-3: intervention at t=0, 80, 140. Column 4: original audio through VAE.
@@ 4.1 Original Audio Reconstruction
To edit a real sample, text prompts serve as critical conditions to guide the sampling process, requiring a balance between identity preservation in latents and responsiveness to prompt guidance. Before fusion, we first examine the resistance of latent codes produced by our inversion models to prompt conditioning. We use null prompt embeddings in all experiments in this subsection.
@@ 4.1.1 Necessity of Omitting Stochastic Terms
Stochastic terms are added back during sampling because they help preserve the potential for editing. Let $\tilde{x}_{t}$ be given by Eq. 6; then the comparisons are as follows:
@@
The first observation is illustrated in Figure 2: any form of noise added during inversion, including retaining stochastic terms or adding noise as in forward diffusion, results in a very noisy sampling phase. This observation demonstrates that omitting noise in inversion is reasonable, and that approximating inversion with diffusion produces samples of unacceptable quality with DPMSolver++. The original audio is reconstructed through VAE encoding and decoding to eliminate the influence of the VAE itself.
4.1.2 Potential of Initial Latent
Figure 3: Comparison of audio generated from initial latents under different methods. Prompt: ”An airplane takes off.” (a) Original audio through VAE. (b) Sample from scratch. (c) Our inversion approach. (d) Audio-to-audio adaptation on Stable Audio Open 1.0. (e) Built-in audio-to-audio of Stable Audio 2.0.
We evaluate the SDE inversion model independently on the initial latent. This evaluation is based on the assumption that, apart from the initial latent representation, the model, sampling process, parameters, and other conditions remain the same. The accuracy with which the initial latent representation captures the characteristics of the original audio determines whether the interpolations in the generation process can precisely produce sound containing the original audio’s features and the degree of fidelity. We compare our method with existing approaches. The comparison methods include:

### v2

# Jing Liu, Enqi Lian, Moyao Deng

@@
Abstract Sound Fusion with Unconditional Inversion Models
Jing Liu, Enqi Lian, Moyao Deng
Department of Music Artificial Intelligence and Music Information Technology
Central Conservatory of Music
Beijing, China
{jing.liu, enqilian, 21A017}@mail.ccom.edu.cn
Correspondence to Jing Liu.
@@ 1 Introduction
Sound fusion seeks to combine an original sound with a reference sound, creating a new auditory output whose acoustic characteristics transcend simple additive mixing of the input components. In this study, sounds are classified into two primary categories: concrete sounds and abstract sounds. Concrete sounds are those that listeners can clearly associate with specific objects, scenes, events, or musical patterns that are intuitively recognizable due to their semantic content. Examples such as running water, crying children, arriving trains, and metallic percussion can be readily identified as concrete. Conversely, abstract sounds are those that do not reveal identifiable real-world sound events or, in musical contexts—particularly in contemporary, electronic, and experimental music clips—do not intuitively disclose recognizable patterns. In essence, the presence of describable, semantically meaningful patterns determines whether a sound is classified as concrete or abstract.
@@
This research specifically focuses on the fusion of abstract sounds; such fused abstract sounds demonstrate musical potential that serve as an emergent material for modern composition. We use "sound" to distinguish our focus from music, and "audio" as the carrier medium of sound. The main contributions of this work are as follows:
@@
  • We introduce an audio fusion approach that synthesizes auditory features from an original audio and a reference audio.
  • We propose both Stochastic Differential Equation (SDE) and Ordinary Differential Equation (ODE) inversion models based on DPMSolver++ samplers that enable this audio fusion capability.
@@ 2 Related Works
Recent research in real audio editing has drawn inspiration from image editing approaches null-text ; edit-friendly ; pivotal_tuning ; SDEdit ; sto_diff_editing , introducing new editing tasks such as substituting sound events within audio, as well as adding, removing, replacing, and refining concrete audio content AUDIT ; PPAE . This research scope expands from concrete audio to encompass abstract audio.
@@
Most pre-existing inversion studies have aimed to reverse samples back to their initial latent codes $x_{T}$, from which sampling begins. This technique originated during the GAN era manifold ; GAN_inversion_survey and has since been extended to diffusion models EDICT ; BELM . From an editing perspective, inversion addresses the challenges of unexpected changes and unwanted feature preservation that occur when employing sample-to-sample approaches.
@@
Prior audio inversion approaches have worked with Denoising Diffusion Probabilistic Models (DDPM) DDPM and Denoising Diffusion Implicit Models (DDIM) DDIM , either approximating inversion through diffusion itself or by reversing the sampling process. Huberman-Spiegelglas et al. edit-friendly computed latent codes at subsequent timesteps directly from $x_{0}$, and recent work zero-shot extended this approach to the audio domain. Mokady et al. null-text worked with the DDIM sampling ODE and swapped timesteps for inversion approximation. Within the Classifier-Free Guidance (CFG) framework CFG , the authors optimized the negative prompt during inversion while preserving the flexibility of the positive prompt. This approach fine-tuned interpolations against diffusion latents as pivots pivotal_tuning . However, fine-tuning embeddings for each edit incurs relatively high computational costs. PPAE PPAE subsequently utilized this approach for audio editing.
We choose the open-source Text-To-Audio (TTA) system Stable Audio Open 1.0 stable_audio_open to achieve the best possible perceptual quality. This system uses DPMSolver++ DPMSolver++ samplers, we introduce DPMSolver++ ODE and SDE inversion models in this work. Since our approach obviates the need for prompt optimization to constrain the inversion trajectory, it preserves prompt flexibility for future creative applications. In this paper specifically, the fusion phase comprises two distinct stages, preceded by fine-tuning the model with the prompt paired with the reference audio. During the initial fusion stage, the prompt effectively guides the sampling trajectory toward reconstructing the reference audio; subsequently, following the intervention of the inverted latent of the original audio, we employ a null text prompt to complete the fusion process.
@@ 3 Methods
In this section, we first introduce the notation employed throughout this paper. Next, we propose inversion models that reconstruct the original audio and describe the fine-tuning of Stable Audio Open to reconstruct reference audio. Finally, we present the audio fusion algorithm.
@@ 3.1 Notation and Definitions
In DDIM and DDPM, $\alpha$ represents the portion of the signal, and $\bar{\alpha}$ is the cumulative product of $\alpha$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$. Although the notation for $\alpha$ and $\bar{\alpha}$ varies in the original sources, we use them consistently throughout this work.
[…]
@@ 3.2 Model Output as Constant
DPMSolver++ introduced model output to replace the predicted noise term, which provides us with a sampling perspective: at every step toward the previous timestep, we desire the noise prediction network to predict a noise map that perfectly matches the latent such that, given both conditions, the model output exactly equals our target sample for inversion. An SDE sampler takes the following form:
$\boldsymbol{x}_{t}=\frac{\sigma_{t}}{\sigma_{s}}e^{-(\lambda_{t}-\lambda_{s})}\boldsymbol{x}_{s}+2\alpha_{t}\int_{\lambda_{s}}^{\lambda_{t}}e^{-2(\lambda_{t}-\lambda)}x_{\theta}^{(\lambda)}\text{d}\lambda+\sqrt{2}\sigma_{t}\int_{\lambda_{s}}^{\lambda_{t}}e^{-(\lambda_{t}-\lambda)}\text{d}\omega_{\lambda}$ (2)
Hence, by assigning the encoded $\boldsymbol{x}_{0}$ of the real sample to $x_{\theta}$ in Eq. 2 and removing the stochastic term, the latent at the next timestep $\boldsymbol{x}_{t}$ is given by:
@@ 3.3 Audio Fusion Algorithm
The fusion process is a modified sampling process. The algorithm operates in two phases: in the first half of fusion, the reference audio prompt dominates to produce interpolations with implicit features of the reference audio. In the second half, we intervene the inverted latent into the sampler.
Each term holds explicit significance at the moment of intervention. As shown in Eq. 2, the first term is where the inverted interpolation is injected, while the second term holds the interpolations guided by the reference prompt. Collectively, these two terms represent an addition of two audio sources, while the second and third terms provide flexibility for fusion.
@@ 4 Experiments and Results
Perceptual similarity is our primary concern for sound fusion. For general audio, especially abstract audio defined in our work, there is currently a lack of valid objective perceptual measurements due to unclear semantics, spectral complexity, and psychoacoustic influences. There exist perceptual metrics for speech JND ; however, speech occupies a narrow frequency range, whereas general audio spans the full spectrum with diverse timbral qualities, complex textures, and non-harmonic elements that speech metrics cannot adequately capture. Therefore, we employ spectrograms for intuitive visual analysis, presenting fusion results and subjective evaluation to assess fusion effects, while we assess signal quality for reconstruction.
@@ 4.1 Original Audio Reconstruction
We first examine the capability of our inversion models for reconstructing original audio.
@@ 4.1.1 Necessity of Omitting Stochastic Terms
Stochastic terms are added back during sampling because they help preserve the potential for editing. Let $\tilde{x}_{t}$ be given by Eq. 3. We evaluate three different formulations for $x_{t}$: (a) our proposed inversion model, (b) an approximation of prior work edit-friendly adapted to our case, and (c) preserving the stochastic term in our inversion model.
@@
The results illustrate that our approach achieves the best performance among the three methods, where the dashed lines represent the mean performance for each approach in Figure 1. Any form of noise added during inversion—whether adding noise as in forward diffusion or retaining stochastic terms—results in a very noisy sampling phase. Notably, audio clips were reconstructed through Variational AutoEncoder (VAE) encoding and decoding to eliminate the influence of the VAE itself.
Figure 1: Audio Quality Metrics Comparison: Three Inversion Approaches.
4.1.2 Comparisons with Audio-to-Audio Approaches
We compare our method with existing approaches, including: